# VIVAT: Virtuous Improving VAE Training through Artifact Mitigation

### v1

@@ 1 Introduction
Variational Autoencoders (VAEs) [22, 38] have long been established as a cornerstone of generative AI, playing a pivotal role in tasks such as image reconstruction and image generation [53, 37, 47, 13, 55]. Their significance is further emphasized by their integration into modern latent diffusion pipelines [39, 8, 12, 26, 33, 2, 50, 7], where they serve as the main ingredient for latent space formation. Despite their widespread adoption and extensive research history, the training of high-quality VAEs remains fraught with challenges, particularly with respect to the emergence of artifacts that degrade generation quality [25, 43]. These persistent issues underscore the continued relevance of refining VAE training methodologies to enhance their robustness and generative fidelity.
@@
Previous research has explored various architectural modifications and different concepts of VAEs, leading to the development of specialized model families and various changes in the optimization procedure [18, 48, 37, 40, 47, 9, 19, 31, 7, 25]. Although these contributions have expanded the theoretical understanding of VAEs, many works focus on fundamental innovations while neglecting practical considerations, particularly the systematic analysis and mitigation of training artifacts. Consequently, despite the diversity of proposed solutions, there remains a lack of straightforward and actionable strategies to address anomalies in existing VAE implementations and training process. In this work, we conduct a comprehensive classification and analysis of these artifacts, offering practical, easy-to-implement recommendations to eliminate them without resorting to radical architectural overhauls.
@@
We posit that the potential of conventional autoencoding frameworks remains underutilized and that significant improvements can be achieved through targeted refinements rather than entirely new methodologies. To this end, we adopt the classic KL-VAE  [^1] Later in this paper, the terms “VAE” and “KL-VAE” will be used interchangeably, unless otherwise stated. architecture [22] – a proven and widely used framework that has been successfully applied in prior work for modern generative models [8, 26, 12]. By systematically investigating the origins of training artifacts within this architecture, we propose simple yet effective solutions that yield measurable improvements in model performance. Quantitative evaluations demonstrate that our approach achieves state-of-the-art results for reconstruction quality without introducing unnecessary complexity or departing from established VAE training principles. With this work, we aim to provide valuable insights for researchers and practitioners who are seeking to optimize VAE training while preserving the simplicity and efficiency of classical autoencoding methods.
@@
  • Finally, we show that VAE, trained according to our methodology, is well-suited for the text-to-image generation task and demonstrates a high-level quality that surpasses the previous state-of-the-art image reconstruction model, namely Flux VAE [26], in terms of CLIP score.
@@ 2.1 VAE and its modifications
The classical VAE, also referred to as KL-VAE, was introduced by [22] and quickly adopted for conditional image generation [53]. The core principle of this approach involves minimizing the Kullback-Leibler (KL) divergence between the latent distribution and a Gaussian prior. Since its inception, this architecture has been widely utilized and extensively modified. Notable architectural innovations include hierarchical VAE [47] and DiffuseVAE [31], the latter employing diffusion processes to refine the blurry VAE outputs. Previous research has also explored the impact of architectural depth in autoencoders [9] and proposed a VAE variant optimized for high compression ratios (up to 128$\times$) [7]. However, the majority of prior work has focused on developing novel loss functions and regularization techniques, such as spectral regularization [47], frequency-domain optimization [19], the integration of Wasserstein distance [45], and more [18, 40, 25, 43, 11, 44, 42, 57]. Despite these advancements, most of these contributions have not been widely adopted in large-scale commercial generative models. An alternative approach, VQ-VAE [48], replaces the continuous latent space with a discrete codebook, leading to significant improvements in generative tasks [37, 13, 55, 2]. This method has inspired numerous subsequent modifications [37, 58], including VQ-GAN [13], which enhances VQ-VAE through adversarial and perceptual losses. Nevertheless, quantized VAEs are known to produce geometric artifacts, and scaling their latent space dimensions is non-trivial. As a result, KL-VAE remains prevalent in several recent high-profile generative models [8, 12, 26]. In this work, we further investigate the KL-VAE architecture, aiming to advance its practical potential in reconstruction and generative applications.
@@ 2.2 VAE for latent diffusion
Due to inherent limitations in sample quality, VAEs are not the best option for generation tasks on their own. However, they have become a prominent component in latent diffusion models [39]. By reducing the spatial dimensions of images, VAEs enable more efficient representations, replacing computationally expensive pixel-based approaches. Latent diffusion models for image generation employ both continuous KL-VAEs [8, 12, 26] and quantized autoencoders [36, 2]. Key principles include increasing the channel size of latent variables and spatial compression ratios. When extending diffusion pipelines to video generation, modified VAEs were developed to incorporate temporal dependencies, either by preserving channel dimensions [6, 3] or compressing along the temporal axis [54, 24]. Since video VAEs build upon the image VAEs, our work focuses on analyzing spatial artifacts in image reconstruction, without loss of generality for the proposed solutions.
@@ 2.3 VAE artifacts
The problem of artifacts in image reconstruction using VAEs has been widely studied. A well-known limitation of VAEs is their tendency to produce blurred outputs, attributed to the asymmetric nature of the KL divergence [23]. While adversarial loss can mitigate blurring, it often introduces visual artifacts [27] and training instability, particularly at high resolutions [7]. Additionally, VAEs are prone to losing high-frequency information [19]. Stronger regularization (e.g., increasing the KL-divergence weight) yields a smoother latent space but reduces its information capacity, leading to loss of fine details and degraded reconstruction quality [25, 46]. Numerous studies have proposed improved training methods to reduce artifacts, yet most lack a systematic analysis or taxonomy of the specific visual artifacts encountered [40, 13, 19, 31]. [25] highlights a connection between artifacts and insufficient equivariance in the latent space under spatial transformations, though without explicitly categorizing the artifacts. Recently, visual artifacts in VAEs outputs have also been discussed in the context of diffusion-based generation pipelines. [43] notes that high-frequency components in VAE latent spaces can propagate undesirable artifacts into the final RGB images. In this work, we systematize common and problematic artifacts in VAE reconstructions and propose simple, effective mitigation strategies that require neither extensive changes to the training process nor fundamental modifications of the VAE architecture.
@@ 3.1 VAE principles and objectives
The Variational Autoencoder (VAE) [22] is a generative model that learns latent representations of observed data $\mathbf{x}$ through an encoder-decoder architecture. The encoder network parameterizes the approximate posterior distribution $q_{\phi}(\mathbf{z}|\mathbf{x})$ with weights $\phi$, mapping input data to a latent space typically assumed Gaussian, while the decoder network parameterizes the likelihood $p_{\theta}(\mathbf{x}|\mathbf{z})$ with weights $\theta$, reconstructing data from latent samples $\mathbf{z}$. The prior over the latent space $p(\mathbf{z})$ is usually chosen as a standard normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. The VAE optimizes the evidence lower bound (ELBO), which balances reconstruction accuracy and latent space regularization:
@@
KL Divergence Loss [22].
@@
Reconstruction Loss [22].
@@
Adversarial Loss [27].
@@
$\mathcal{L}_{\text{adv}}=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\log(1-D(\hat{\mathbf{x}})),$ (5)
@@
Perceptual Loss [20].
@@
A popular choice for the perceptual loss is LPIPS [56].
@@ 3.2 VAE architecture
Our base VAE architecture builds upon the open-source  [^2] https://github.com/ai-forever/MoVQGAN SBER-MoVQGAN model [2], which itself is derived from VQGAN [13] and incorporates spatially conditional normalization from MoVQ [58]. In our implementation, we replace the quantization bottleneck with mean $\mu$ and log-variance $\log\sigma^{2}$ prediction layers to form a KL-VAE continuous latent space.
Figure 1 provides a detailed scheme of the KL-VAE structures. Following the design in [39], each encoder block consists of a ResNet Block [16] followed by a downscale module, with some layers additionally incorporating Self-Attention Blocks [49]. The decoder similarly employs ResNet blocks but replaces downscale modules with upscale operations, retaining self-attention where applicable. Detailed ResNet and Self-Attention block schemes are presented in Appendix A.1.
@@ 3.3 Latent diffusion VAE
Latent diffusion models (LDMs) [39] perform diffusion processes in a compressed latent space rather than operating directly on high-dimensional input data (e.g., pixel space for images). By employing a variational autoencoder (VAE) to encode data into a lower-dimensional representation, LDMs achieve significant computational efficiency while maintaining high generation quality.
The VAE encoder transforms an input image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ (with height $H$, width $W$, and channels $C$) into a latent representation $\mathbf{z}\in\mathbb{R}^{H/f\times W/f\times c}$, where $f$ is the spatial downscaling factor and $c$ denotes the increased channel dimension for information preservation. In transformer-based implementations [32], the quadratic complexity of full self-attention with respect to latents size makes spatial dimension operations particularly costly. Channel dimension adjustments, however, can be efficiently implemented through linear projections with minimal computational overhead.
In this work, we further investigate the suitability of our enhanced image reconstruction VAE in a latent diffusion pipeline for text-to-image generation. We employ a diffusion transformer based on LI-DiT [29], training it with our frozen VAE, and present the results in Section 6.3.
@@ 4 Image VAE challenges
This peculiarity was also mentioned in the [34]. The authors pointed out that this problem caused by the high norms in latent codes in certain spatial locations. After decoding these high norms produce the bright spots in the pixel space. In our work we also faced with this problem, the example is presented at the Figure 2(f). Also the Figure 3 shows formation process of this artifact by displaying the activations norms for the different layers of the VAE decoder. To mitigate this problem the authors of the [34] invented an additional outlier loss, which penalizes the model for latent values far from the mean. We propose to solve this problem in other, simpler way.
@@ 5.1 Data preprocessing
The resizing method also significantly impacts performance. Our experiments indicate that the default nearest-neighbor method yields suboptimal results, whereas the bicubic method produces superior outcomes. These findings align with established results from prior work [15, 5]. Notably, training a VAE on lower-resolution images (e.g., 240p) can remain effective for operation at higher resolutions (e.g., 480p or 720p), as the convolutional receptive field remains unchanged. Thus, the VAE learns to compress the image by a factor of $n$ along each side in the same manner as with smaller resolutions.
Our complete training dataset consists of 100 million images sourced from the publicly available LAION Aesthetic HighRes dataset, a subset of LAION-5B [41]. All images in this dataset have HD resolution and an associated aesthetic score. We selected only images with a score $>4.5$, resized them from HD to 480p, and then cropped them to 240p.
@@ 5.2 Challenges solutions
At first this artifact appeared in the corner and we assumed that it also appeared due to incorrect padding and that we would get rid of it with the previous artifact after padding changing. But after that the dot began to move throughout the whole picture. As it was already mentioned the real reason is the high norms in activations during decoding. Unlike the [34] we decided not to add the new loss term, as it would make the total loss more complicated and hard to adjust its weights, but make changes to the architecture Of the model. We use the method proposed in [58] named Spatially Conditional Normalization. The authors used this normalization to modulate the quantized vectors so as to insert spatially variant information to the embedded index maps. In this way, they achieved a reduction in the number of artifacts that arise due to encoding similar patches into one latent codebook vector. For us this layer has two main benefits. First, it acts as a supplementary normalization layer. Second, it provides additional spatial interactions between parts of the image. And although we are not working with a discrete latent space, using this layer helps us get rid of the droplet artifact.
@@ 5.3 Decoder finetuning
Another sufficient step to improve VAE reconstruction quality is a freezing of encoder. This training stage can further improve quality of trained model and help to get rid of other artifacts. In our experiments, an additional training phase with a fixed encoder helped us further increase the FID metric at a late phase of training. Similar strategy was used for example in [14] where the stage with frozen encoder was even longer then full VAE training stage.
@@ 6.2 Image reconstruction
To assess the reconstruction performance of our approach, we employ two standard metrics: peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM). We evaluate these metrics on multiple benchmark datasets, including ImageNet [^3] https://www.image-net.org/challenges/LSVRC/2012/ [10] at resolutions of $256\times 256$ and $512\times 512$, MS COCO 2017 Test [^4] https://cocodataset.org/ [28] at $512\times 512$, and FFHQ [^5] https://github.com/NVlabs/ffhq-dataset [21] at $1024\times 1024$.
For comparison, we test our model against several state-of-the-art autoencoders: DC-VAE [7], Cosmos-0.1-Tokenizer-CI8$\times$8666https://huggingface.co/nvidia/Cosmos-0.1-Tokenizer-CI8x8 [30], Flux VAE [^7] https://huggingface.co/black-forest-labs/FLUX.1-dev [26], and the VAE from Stable Diffusion 3 [12]. The quantitative results, presented in Table 1, demonstrate that our method achieves state-of-the-art performance in image reconstruction, attaining the highest – or competitively equivalent – scores in both SSIM and PSNR across all evaluated datasets. This quantitative superiority is further supported by qualitative evaluation, as illustrated in Figure 5, where our method exhibits competitive results compared to the best approaches.

### v2

@@ 1 Introduction
Variational Autoencoders (VAEs) [Kingma2013AutoEncodingVB, pmlr-v32-rezende14] have long been established as a cornerstone of generative AI, playing a pivotal role in tasks such as image reconstruction and image generation [10.1007/978-3-319-46493-0_47, 10.5555/3454287.3455618, vahdat2020NVAE, esser2020taming, yu2022vectorquantizedimagemodelingimproved]. Their significance is further emphasized by their integration into modern latent diffusion pipelines [Rombach2021HighResolutionIS, chen2023pixartalpha, esser2024scalingrectifiedflowtransformers, flux2024, podell2023sdxlimprovinglatentdiffusion, arkhipkin2024kandinsky30technicalreport, vladimir-etal-2024-kandinsky, chen2024deep], where they serve as the main ingredient for latent space formation. Despite their widespread adoption and extensive research history, the training of high-quality VAEs remains fraught with challenges, particularly with respect to the emergence of artifacts that degrade generation quality [kouzelis2025eqvaeequivarianceregularizedlatent, skorokhodov2025improvingdiffusabilityautoencoders]. These persistent issues underscore the continued relevance of refining VAE training methodologies to enhance their robustness and generative fidelity.
@@
Previous research has explored various architectural modifications and different concepts of VAEs, leading to the development of specialized model families and various changes in the optimization procedure [Higgins2016betaVAELB, 10.5555/3295222.3295378, 10.5555/3454287.3455618, pmlr-v139-rybkin21a, vahdat2020NVAE, child2021deepvaesgeneralizeautoregressive, jiang2021focal, pandey2022diffusevaeefficientcontrollablehighfidelity, chen2024deep, kouzelis2025eqvaeequivarianceregularizedlatent]. Although these contributions have expanded the theoretical understanding of VAEs, many works focus on fundamental innovations while neglecting practical considerations, particularly the systematic analysis and mitigation of training artifacts. Consequently, despite the diversity of proposed solutions, there remains a lack of straightforward and actionable strategies to address anomalies in existing VAE implementations and training process. In this work, we conduct a comprehensive classification and analysis of these artifacts, offering practical, easy-to-implement recommendations to eliminate them without resorting to radical architectural overhauls.
@@
We posit that the potential of conventional autoencoding frameworks remains underutilized and that significant improvements can be achieved through targeted refinements rather than entirely new methodologies. To this end, we adopt the classic KL-VAE  [^1] Later in this paper, the terms “VAE” and “KL-VAE” will be used interchangeably, unless otherwise stated. architecture [Kingma2013AutoEncodingVB] – a proven and widely used framework that has been successfully applied in prior work for modern generative models [chen2023pixartalpha, flux2024, esser2024scalingrectifiedflowtransformers]. By systematically investigating the origins of training artifacts within this architecture, we propose simple yet effective solutions that yield measurable improvements in model performance. Quantitative evaluations demonstrate that our approach achieves state-of-the-art results for reconstruction quality without introducing unnecessary complexity or departing from established VAE training principles. With this work, we aim to provide valuable insights for researchers and practitioners who are seeking to optimize VAE training while preserving the simplicity and efficiency of classical autoencoding methods.
@@
  • Finally, we show that VAE, trained according to our methodology, is well-suited for the text-to-image generation task and demonstrates a high-level quality that surpasses the previous state-of-the-art image reconstruction model, namely Flux VAE [flux2024], in terms of CLIP score.
@@ 2.1 VAE and its modifications
[…]
@@ 2.2 VAE for latent diffusion
Due to inherent limitations in sample quality, VAEs are not the best option for generation tasks on their own. However, they have become a prominent component in latent diffusion models [Rombach2021HighResolutionIS]. By reducing the spatial dimensions of images, VAEs enable more efficient representations, replacing computationally expensive pixel-based approaches. Latent diffusion models for image generation employ both continuous KL-VAEs [chen2023pixartalpha, esser2024scalingrectifiedflowtransformers, flux2024] and quantized autoencoders [ramesh2022hierarchicaltextconditionalimagegeneration, arkhipkin2024kandinsky30technicalreport]. Key principles include increasing the channel size of latent variables and spatial compression ratios. When extending diffusion pipelines to video generation, modified VAEs were developed to incorporate temporal dependencies, either by preserving channel dimensions [blattmann2023videoldm, 10815947] or compressing along the temporal axis [yang2025cogvideoxtexttovideodiffusionmodels, kong2025hunyuanvideosystematicframeworklarge]. Since video VAEs build upon the image VAEs, our work focuses on analyzing spatial artifacts in image reconstruction, without loss of generality for the proposed solutions.
@@ 2.3 VAE artifacts
The problem of artifacts in image reconstruction using VAEs has been widely studied. A well-known limitation of VAEs is their tendency to produce blurred outputs, attributed to the asymmetric nature of the KL divergence [Kingma_2019]. While adversarial loss can mitigate blurring, it often introduces visual artifacts [Larsen2015AutoencodingBP] and training instability, particularly at high resolutions [chen2024deep]. Additionally, VAEs are prone to losing high-frequency information [jiang2021focal]. Stronger regularization (e.g., increasing the KL-divergence weight) yields a smoother latent space but reduces its information capacity, leading to loss of fine details and degraded reconstruction quality [kouzelis2025eqvaeequivarianceregularizedlatent, 10.1007/978-3-031-72998-0_17]. Numerous studies have proposed improved training methods to reduce artifacts, yet most lack a systematic analysis or taxonomy of the specific visual artifacts encountered [pmlr-v139-rybkin21a, esser2020taming, jiang2021focal, pandey2022diffusevaeefficientcontrollablehighfidelity]. [kouzelis2025eqvaeequivarianceregularizedlatent] highlights a connection between artifacts and insufficient equivariance in the latent space under spatial transformations, though without explicitly categorizing the artifacts. Recently, visual artifacts in VAEs outputs have also been discussed in the context of diffusion-based generation pipelines. [skorokhodov2025improvingdiffusabilityautoencoders] notes that high-frequency components in VAE latent spaces can propagate undesirable artifacts into the final RGB images. In this work, we systematize common and problematic artifacts in VAE reconstructions and propose simple, effective mitigation strategies that require neither extensive changes to the training process nor fundamental modifications of the VAE architecture.
@@ 3.1 VAE principles and objectives
The Variational Autoencoder (VAE) [Kingma2013AutoEncodingVB] is a generative model that learns latent representations of observed data $\mathbf{x}$ through an encoder-decoder architecture. The encoder network parameterizes the approximate posterior distribution $q_{\phi}(\mathbf{z}|\mathbf{x})$ with weights $\phi$, mapping input data to a latent space typically assumed Gaussian, while the decoder network parameterizes the likelihood $p_{\theta}(\mathbf{x}|\mathbf{z})$ with weights $\theta$, reconstructing data from latent samples $\mathbf{z}$. The prior over the latent space $p(\mathbf{z})$ is usually chosen as a standard normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. The VAE optimizes the evidence lower bound (ELBO), which balances reconstruction accuracy and latent space regularization:
@@
KL Divergence Loss [Kingma2013AutoEncodingVB].
@@
Reconstruction Loss [Kingma2013AutoEncodingVB].
@@
Adversarial Loss [Larsen2015AutoencodingBP].
@@
$),\mathcal{L}_{\text{adv}}=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\log(1-D(\hat{\mathbf{x}}$)), (5)
@@
Perceptual Loss [Johnson2016PerceptualLF].
@@
A popular choice for the perceptual loss is LPIPS [Zhang2018TheUE].
@@ 3.2 VAE architecture
Our base VAE architecture builds upon the open-source  [^2] https://github.com/ai-forever/MoVQGAN SBER-MoVQGAN model [arkhipkin2024kandinsky30technicalreport], which itself is derived from VQGAN [esser2020taming] and incorporates spatially conditional normalization from MoVQ [zheng2022movq]. In our implementation, we replace the quantization bottleneck with mean $\mu$ and log-variance $\log\sigma^{2}$ prediction layers to form a KL-VAE continuous latent space.
Figure 1 provides a detailed scheme of the KL-VAE structures. Following the design in [Rombach2021HighResolutionIS], each encoder block consists of a ResNet Block [He2015DeepRL] followed by a downscale module, with some layers additionally incorporating Self-Attention Blocks [Vaswani2017AttentionIA]. The decoder similarly employs ResNet blocks but replaces downscale modules with upscale operations, retaining self-attention where applicable. Detailed ResNet and Self-Attention block schemes are presented in Appendix A.1.
@@ 3.3 Latent diffusion VAE
Latent diffusion models (LDMs) [Rombach2021HighResolutionIS] perform diffusion processes in a compressed latent space rather than operating directly on high-dimensional input data (e.g., pixel space for images). By employing a variational autoencoder (VAE) to encode data into a lower-dimensional representation, LDMs achieve significant computational efficiency while maintaining high generation quality.
The VAE encoder transforms an input image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ (with height $H$, width $W$, and channels $C$) into a latent representation $\mathbf{z}\in\mathbb{R}^{H/f\times W/f\times c}$, where $f$ is the spatial downscaling factor and $c$ denotes the increased channel dimension for information preservation. In transformer-based implementations [Peebles2022ScalableDM], the quadratic complexity of full self-attention with respect to latents size makes spatial dimension operations particularly costly. Channel dimension adjustments, however, can be efficiently implemented through linear projections with minimal computational overhead.
In this work, we further investigate the suitability of our enhanced image reconstruction VAE in a latent diffusion pipeline for text-to-image generation. We employ a diffusion transformer based on LI-DiT [ma2024exploring], training it with our frozen VAE, and present the results in Section 6.3.
@@ 4 Image VAE challenges
This peculiarity was also mentioned in the [polyak2024moviegen]. The authors pointed out that this problem caused by the high norms in latent codes in certain spatial locations. After decoding these high norms produce the bright spots in the pixel space. In our work we also faced with this problem, the example is presented at the Figure 2(f). Also the Figure 3 shows formation process of this artifact by displaying the activations norms for the different layers of the VAE decoder. To mitigate this problem the authors of the [polyak2024moviegen] invented an additional outlier loss, which penalizes the model for latent values far from the mean. We propose to solve this problem in other, simpler way.
@@ 5.1 Data preprocessing
The resizing method also significantly impacts performance. Our experiments indicate that the default nearest-neighbor method yields suboptimal results, whereas the bicubic method produces superior outcomes. These findings align with established results from prior work [Han2013, Bhatt]. Notably, training a VAE on lower-resolution images (e.g., 240p) can remain effective for operation at higher resolutions (e.g., 480p or 720p), as the convolutional receptive field remains unchanged. Thus, the VAE learns to compress the image by a factor of $n$ along each side in the same manner as with smaller resolutions.
Our complete training dataset consists of 100 million images sourced from the publicly available LAION Aesthetic HighRes dataset, a subset of LAION-5B [schuhmann2022laion]. All images in this dataset have HD resolution and an associated aesthetic score. We selected only images with a score >4.5>4.5, resized them from HD to 480p, and then cropped them to 240p.
@@ 5.2 Challenges solutions
At first this artifact appeared in the corner and we assumed that it also appeared due to incorrect padding and that we would get rid of it with the previous artifact after padding changing. But after that the dot began to move throughout the whole picture. As it was already mentioned the real reason is the high norms in activations during decoding. Unlike the [polyak2024moviegen] we decided not to add the new loss term, as it would make the total loss more complicated and hard to adjust its weights, but make changes to the architecture Of the model. We use the method proposed in [zheng2022movq] named Spatially Conditional Normalization. The authors used this normalization to modulate the quantized vectors so as to insert spatially variant information to the embedded index maps. In this way, they achieved a reduction in the number of artifacts that arise due to encoding similar patches into one latent codebook vector. For us this layer has two main benefits. First, it acts as a supplementary normalization layer. Second, it provides additional spatial interactions between parts of the image. And although we are not working with a discrete latent space, using this layer helps us get rid of the droplet artifact.
@@ 5.3 Decoder finetuning
Another sufficient step to improve VAE reconstruction quality is a freezing of encoder. This training stage can further improve quality of trained model and help to get rid of other artifacts. In our experiments, an additional training phase with a fixed encoder helped us further increase the FID metric at a late phase of training. Similar strategy was used for example in [audiovae] where the stage with frozen encoder was even longer then full VAE training stage.
@@ 6.2 Image reconstruction
To assess the reconstruction performance of our approach, we employ two standard metrics: peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM). We evaluate these metrics on multiple benchmark datasets, including ImageNet [^3] https://www.image-net.org/challenges/LSVRC/2012/ [5206848] at resolutions of $256\times 256$ and $512\times 512$, MS COCO 2017 Test [^4] https://cocodataset.org/ [lin2015microsoftcococommonobjects] at $512\times 512$, and FFHQ [^5] https://github.com/NVlabs/ffhq-dataset [karras2019stylebasedgeneratorarchitecturegenerative] at $1024\times 1024$.
[…]